This is a report on the collection activities from the past 4 days.
Thursday 7/24: I was surprised to see a sample this morning, a little over 100 ml. When I arrived at 8:30 the doors were open and the fan was running yet the fog had ended. I rinsed the system, but closed it down from the fogPI, since I did not want it to be open all day. I knew someone would come later to take a blank and they could restart it then. Chris and Erin came at 1 pm and took a blank, cycled power, and put on a bottle.
As an aside, a spider trap was placed a day or 2 before.
Friday 7/25: When I drove by the system that morning, expecting a sample, it was closed up. Apparently, cycling the power was insufficient to restart fogPI. I came by later that afternoon and found that the bottle had also fallen off. I reinitiated the fogPI through the system and installed a new bottle.
Saturday 7/26: No sample, though we thought there had been fog. It had opened. Chris, Erin, Cristian and I acid washed the system.
Sunday 7/27: The bottle had fallen off again! This was the 3rd time! We need to tighten it fully and then just slightly loosen it to allow air to flow out, but not loosen it more than just that slight amount due to the conditions of the site. Since it was opened and bugs had flown in, I carefully cleaned the strings by pulling them out and rinsing them with DI water. With a clean hand, I reinstalled them.
No spiders yet.